Grade 12Algebra

Let f(x)=x^2+b1x+c1, g(x)=x^2+b2x+c2. Real roots of f( x) be A,B and real roots of g(x) be A+D and B+D. Least value of f(x) be -1/4. Least value of g(x) occurs at x=7/2. Find least value of g(z).

Tyler
8 Years agoGrade 12

1 Answer

Arun
8 Years ago
Dear student
 
least value of f(x) =  – (b12 – 4c1)/4 = – ¼
Hence
(b12 – 4c1) = 1
Now
 – b2 / 2 = 7/2
 – b2  = – 7
Now find roots of F(x) by formula and you will get those two roots.